Night Center: 00:33

Handover: 00:23 (take into account extreme changes between focal stations)

ISA: D. Thompson (Remote Room)
  • Complete LUCI-2/AGW-2 On-Sky Checks - high priority - J. Hill/D. Thompson - 1-2 hours
    • See UT-2020-01-30 Wiki page for gory details
    • Files located on:
      • RM507-5/4: /home/observer/share/Service_Observing/Eng_Jan_2020/LUCI2_AGW_Checkout
      • OBS-N: /home/lbto/Service_Observing/Eng_Jan_2020/LUCI2_AGW_Checkout
    • Verify daytime cable fix did not incur other issues - check off-axis guiding at various distances in the patrol field
    • Verify S-H spots are centered - adjust if needed
    • Measure Center of Rotation - adjust if needed
    • LUCI Detector Rotator Center (2 opposing PAs, 2 pairs if time)
    • Verify Rotator ZP with Astrometric Masks - tolerates up to 1.2" seeing
      • INAF mask had 3 stars - only 1 was properly centered, updated mask was done wrong and did not work on-sky (PI used wrong version of LMS)
      • Rotator offset gave two different answers depending on machine used
        • RM507-5: Angle = -0.0479 err RA =0.073 err Dec = 0.123
        • OBS-4: Angle = -0.00072 Err RA = 0.073 and Err Dec = 0.134
        • Amended mask run by D. Thompson Angle = 0.09366, err RA=0.076 err Dec=0.051 - done offline UT 03-Feb-2020
        • Current Rotator ZP is 177.05 (this value was confirmed as the one used in 2019B - large offsets seemed to work for telluric)
    • Take Confirmatory Transform Data - Stone Field - up to 1.5" seeing on DIMM
      • Send an Active preset to Stone D or Stone E, collimate
      • Then run transform_collect as in instructions
      • Config for LUCI-2
        • N30 FieldStop
        • 2.51sec DIT
        • H-band or FeII (take a test exposure in the field)
      • lists are located here on OBS-N: /home/lbto/idl/wfsc/transform/lists
      • use:
        • Stone_D_LUCI_PA0_short.list
        • Stone_Em135_0.list ( don't have to complete full list)
  • AO / IRTC Work - AO on-sky data collection tasks - pending release of IRTC (computer issues) and confirmation from AO group they need/want time - IN Parallel with LUCI-2/AGW-2 work above - D. Miller
  • Switch to MODS for D-time - Rollover from Previous Night Plans
    • Files are located here: /home/lbto/Service_Observing/Eng_Jan_2020
      • Each PI has their own sub-folder, each folder has a README and finder charts
      • All scripts have been generated from the OT, but original OT files are in each directory
      • There is a common MODS calibrations folder - includes the Spectro-photometric Standard G191B2b - up for all of first half
      • There are 4 pdf files that show visibility for the targets over the next two nights
    • Spectrophotometric Standard: G191b2b - Dual Grating Only
      • If IRTC work cannot be done:
        • M33-7 - MODS-Bino visible until 10pm - M. Edwards
          • 01 33 35.50 +30 39 29.0
          • if seeing < 0.8" 1.5 hours (thin cirrus OK) - 0.8" slit
          • if seeing > 0.8" 2.5 hours (thin cirrus OK) - 1.0" slit
      • Once IRTC attempt is completed - OR - if IRTC not done (continue from M33-7):
        • V392 - visible through handover - M. Wagner (preferred until 9:30)
          • 06:09:57.453 12:12:25.15
          • Seeing < 1.5" - thin cirrus acceptable
          • 0.8" slit
          • 0.75 hour
        • NGC 2403-5 - visible through handover - M. Edwards
          • 07 36 54.30 +65 36 07.5
          • if seeing < 0.8" 1.5 hours (thin cirrus OK) - 0.8" slit
          • if seeing > 0.8" 2.5 hours (thin cirrus OK) - 1" slit
        • V2860 Ori - visible through handover - M. Wagner
          • 06:09:57.453 12:12:25.15
          • Seeing < 1.5" - thin cirrus acceptable
          • 0.8" slit
          • 0.75 hour
        • V959 Mon- visible through handover M. Wagner
          • 06:39:38.61 05:49:26.74
          • Seeing < 1.5" - thin cirrus acceptable
          • 0.8" slit
          • 0.75 hour
        • V435 CMa - low declination target - 9pm-handover - never rises above 35 elevation - M. Wagner
          • 07:13:45.84 -21:12:31.3
          • Seeing < 1.5" - thin cirrus acceptable
          • 0.8" slit
          • 0.75 hour
  • Poor Weather (seeing > 1.5" and/or more than thin cirrus) - Rollover from Previous Night Plans
    • LUCI-2/LBC-Blue mixed mode tests
      • File Location OBS-N: /home/lbto/Service_Observing/Eng_Jan_2020/Eng_Dec_MixedMode/
      • Readme in directory
    • Local Mergers LUCI-2 zJ/HKspec - B. Rothberg
      • Files are located on OBS-N: /home/lbto/Service_Observing/Eng_Jan_2020/Rothberg_Local_Mergers/
      • Arp 187 - 05:04:53 -10:14:52 - visible until 11:30pm - 2 hours (covers zJ and HKspec observations)
    • Local Mergers - MODS-Bino - B. Rothberg
      • Files are located on OBS-N: /home/lbto/Service_Observing/Eng_Jan_2020/Rothberg_Local_Mergers/
      • UGC 4079 - 07:55:06 +55:42:13 - 1 hour
      • NGC 3310 - 10:38:45 +53:30:11 - 1 hour
      • NGC 2744 - 09:04:39 +18:27:52 - 1 hour
Super Poor Weather
    • H-band Libray - LUCI2 - 30min each - >>2" seeing and/or >2 mags extinction
      • STAR: RA Dec Class Hmag
      • HD 20797 03 24 40.55 +64 35 09.57 M0Iab
      • HD 42543 06 12 19.10 +22 54 30.7 M2I
      • HD 44537 06 24 53.90 +49 17 16.4 M2I
      • BD -15 1319 06 15 01.88 -15 52 49.7 K5III
-- BarryRothberg - 03 Feb 2020
Topic revision: r3 - 04 Feb 2020, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback